OpenStudy (anonymous):

w^2-2w-3/w^2-2w+1 * w-1/w-3 multiply and simplify

OpenStudy (anonymous):

=(w-3)(w+1)/(w-1)^2 *(w-1)/(w-3) cancel out (w-1) & (w-3)

OpenStudy (anonymous):

=(w+1)/(w-1)

OpenStudy (anonymous):

Once you multiply you get (w^3-2w^2-3w-w^2+2w+3)/(w^3-2w^2+w-3w^2+6w-3) Then combine like terms to get (w^3-3w^2-w+3)/(w^3-5w^2+7w-3)

OpenStudy (anonymous):

Pfft inik did it the easier way....why you so good!? XD

OpenStudy (anonymous):

i'm lazy... hate long multiplications :)

OpenStudy (anonymous):

XD My pre-cal teacher does a bunch of the long way ones, he's evil.

OpenStudy (anonymous):

said answer was wrong

OpenStudy (anonymous):

oups! let's check!

OpenStudy (anonymous):

nope... still getting: (w+1)/(w-1) did you check original problem?

OpenStudy (anonymous):

it's (w+1) divide by (w-1)... please check again. I do not see any mistakes

OpenStudy (anonymous):

it wanted w-1/w-3

OpenStudy (anonymous):

how? did you see the solution above? w^2-2w-3=(w-3)(w+1) so, you have to have (w+1) in answer... can't get it. check with your teacher... something is wrong